Ling is cuttinh carpet for a rectangular room. The area of the room is 324 ft^2. The length of the room is 3 feet longer than twice the width. What should the dimensions of the carpet be?
Answer by stanbon(75887)

Ling is cutting carpet for a rectangular room. The area of the room is 324 ft^2. The length of the room is 3 feet longer than twice the width. What should the dimensions of the carpet be?
Width = W
Length = 2W+3
-----
Equation:
W(2W+3) = 324
2W^2 + 3W - 324 = 0
-------
(W -12)(2W+27) = 0
Width = 12 ft
Length = 2W+3 = 27 ft
------------
Cheers,
Stan H.
-----------